# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v1

@@ Abstract.
Voice Authentication Systems (VAS) use unique vocal characteristics for verification. They are increasingly integrated into high-security sectors such as banking and healthcare. Despite their improvements using deep learning, they face severe vulnerabilities from sophisticated threats like deepfakes and adversarial attacks. The emergence of realistic voice cloning complicates detection, as systems struggle to distinguish authentic from synthetic audio. While anti-spoofing countermeasures (CMs) exist to mitigate these risks, many rely on static detection models that can be bypassed by novel adversarial methods, leaving a critical security gap. To demonstrate this vulnerability, we propose the Spectral Masking and Interpolation Attack (SMIA), a novel method that strategically manipulates inaudible frequency regions of AI-generated audio. By altering the voice in imperceptible zones to the human ear, SMIA creates adversarial samples that sound authentic while deceiving CMs. We conducted a comprehensive evaluation of our attack against state-of-the-art (SOTA) models across multiple tasks, under simulated real-world conditions. SMIA achieved a strong attack success rate (ASR) of at least 82% against combined VAS/CM systems, at least 97.5% against standalone speaker verification systems, and 100% against countermeasures. These findings conclusively demonstrate that current security postures are insufficient against adaptive adversarial attacks. This work highlights the urgent need for a paradigm shift toward next-generation defenses that employ dynamic, context-aware frameworks capable of evolving with the threat landscape.
@@
$c$s: Security and privacy Biometricsccs: Computing methodologies Artificial intelligenceccs: Security and privacy Systems security
@@ 1. Introduction
Voice authentication has emerged as a widely adopted biometric technology due to its ease of use and flexibility in enhancing security within various scenarios, including mobile devices, virtual assistants, secure communication systems, financial service institutions, and law enforcement and jurisdictional departments (Singh et al., 2012; Algabri et al., 2017; Rose, 2006). These systems analyze an individual’s voice to create a distinct vocal signature used for identity verification.
The global voice biometrics market was approximately USD 2.5 billion in 2023 and is projected to reach between USD 11.5 billion and USD 15.7 billion by 2032, growing at a compound annual growth rate (CAGR) ranging from 18.7% to 27.5% during this period (Allied Market Research, 2023). This growth is driven by the increased demand for robust fraud detection and prevention systems, particularly within the financial services sector (Markets and Markets, 2023; Straits Research, 2023; Fortune Business Insights, 2024).
@@
Compared to passwords or access cards, voice authentication offers reliable user verification, reduces fraud risk, and simplifies the user experience by eliminating the need to remember multiple credentials and avoiding password-related vulnerabilities. It is also more stable and cost-effective than other biometric systems. Unlike facial recognition, which requires frequent updates due to changes in appearance, voice characteristics remain relatively consistent after adulthood. Additionally, it relies on affordable microphones and requires less storage than high-resolution cameras and facial data. Additionally, it is easier to integrate with cloud services and multiple devices (Renz et al., 2023; Springer, 2023; Techlocity, 2021; Khan and Zahid, 2010).
@@
Recent advancements in deep learning have revolutionized voice authentication systems (VASs), significantly improving accuracy, robustness, and scalability. Modern deep learning-based approaches, such as Deep Neural Network (DNN) models, have enhanced the field by creating digital voice fingerprints that serve as reference points for authentication, for example, Deep Speaker (Li et al., 2017) and X-Vectors (Snyder et al., 2018). Despite these improvements, VAS remain highly vulnerable to various forms of adversarial attacks (Lan et al., 2022).
@@
In parallel, deep learning has also transformed the field of synthetic speech generation, making it easier than ever to replicate someone’s voice with near-human indistinguishability. Technologies like text-to-speech (TTS) (Liao et al., 2024; Qin et al., 2023; Amazon Web Services, 2025) and voice conversion (VC) (Li et al., 2023; Huang et al., 2024) can generate highly realistic speech using minimal input data. These tools, while impressive, have raised significant concerns regarding their misuse in security-sensitive applications, particularly when synthetic speech is used to deceive or bypass Voice Authentication Systems (VAS) (Shonesy, 2015; Alali and Theodorakopoulos, 2025).
@@
To combat these deepfake attacks, various anti-spoofing countermeasures (CMs), like RawNet2 (Tak et al., 2021b) or RawGAT-ST (Tak et al., 2021a), have been developed to distinguish between genuine and synthetic speech by identifying subtle artifacts that differentiate the two. However, attackers have engineered a new class of threat, known as adversarial spoofing attacks, designed specifically to deceive these CMs. This hybrid approach begins by generating a convincing spoofed audio sample using TTS or VC. Subsequently, a carefully crafted and often imperceptible adversarial perturbation is added to this synthetic speech. The resulting signal has a dual objective: to bypass the primary VAS by mimicking the target speaker while simultaneously fooling the CM into classifying the spoofed audio as authentic, posing a compounded threat to security protocols (Hai et al., 2023; Kassis and Hengartner, 2023).
@@
A compelling example is SiFDetectCracker (Hai et al., 2023), an adversarial attack that can deceive state-of-the-art CMs by only modifying speaker-irrelevant features, such as background noise and silent segments. This method achieved an average success rate of over 80% without altering the speaker’s actual voice, demonstrating that detectors often over-rely on secondary acoustic cues. However, the effectiveness of these attacks is significantly reduced in black-box transferability tests, and their impact on VAS or hybrid systems, which consist of VAS and CMs, remains unexplored. This highlights a critical gap: the need for robust attacks that are not only effective but also transferable across different systems in real-world scenarios.
@@
Recent attacks on VAS have shifted focus from fooling speaker verification systems to breaking their CMs. A pivotal work by Kassis et al. (Kassis and Hengartner, 2023) demonstrated the feasibility of a practical, model-agnostic attack that bypasses these CMs not through query-based optimization, but by applying a series of universal signal-processing transformations to spoofed audio. Their approach, which manipulates cues like silence and spectral energy distribution, is capable of bypassing the entire VA stack. However, their results reveal a significant performance drop against more modern and robust defenses; for instance, the success rate fell to just 15.55% against a combined defense using the AWS Voice ID (Amazon, 2025) and RawDarts (Ge et al., 2021). This inconsistency highlights that while targeting universal flaws in CMs is a promising direction, a critical gap remains in developing attacks that can reliably defeat the current generation of state-of-the-art, hardened VA systems. Moreover, designated-app attack vectors often require privileged device modifications, such as rooting a phone to inject audio directly, which poses a substantial barrier for real-world exploitation in security-critical contexts.
Our main contributions of this paper are summarized as follows:
@@
  • We conduct extensive evaluations of SMIA against widely adopted open-source VAS (Deep Speaker (Li et al., 2017) and X-Vectors (Snyder et al., 2018)), commercial platforms (Microsoft Azure SV (Microsoft, [n. d.])), and state-of-the-art CMs (RawNet2 (Tak et al., 2021b), RawGAT-ST (Tak et al., 2021a), and RawPC-Darts (Ge et al., 2021)).
@@
Unlike prior work, which explored TTS-based attacks, targeted voice authentication systems or countermeasures in isolation, or relied on transferability-based attacks with limited effectiveness, we present a black-box attack that simultaneously compromises both. Our evaluations span a wide range of models, tasks, and real-world scenarios, demonstrating the consistency and severity of this vulnerability. The results reveal that VAS are significantly more fragile than previously assumed, underscoring the urgent need for stronger defenses and providing a solid foundation for future research in biometric security.
This paper is organized as follows. Section 2 provides an overview of voice authentication systems and their associated threats, related works, and the threat model. Section 3 outlines our methodology and describes the SMIA workflow. Section 4 details the experimental setup, including the datasets, models, and configurations used. Section 5 presents the evaluation of our attack. Section 6 reports the results of our experiments, highlighting the effectiveness and limitations of the studied approaches. Section 7 provides an ablation study, analyzing the impact of different components and design choices. Section 8 discusses the implications of our findings and covers the challenges and open questions in securing voice authentication systems. Section 9 outlines future work, suggesting directions to improve system robustness and defense mechanisms. Finally, Section 10 concludes the paper by summarizing the main insights and contributions.
@@ 2. Background and Related Works
2.1. List of Acronyms
For clarity, Table 1 defines the acronyms used throughout this document.
Table 1. List of Acronyms
Acronym Definition
ABTN Adversarial Biometrics Transformation Network
ASR Attack Success Rate
CAGR Compound Annual Growth Rate
CM Countermeasure
CQCCs Constant-Q Cepstral Coefficients
CSI Closed-Set Identification
DNN Deep Neural Network
EER Equal Error Rate
FGSM Fast Gradient Sign Method
GAN Generative Adversarial Network
GAT Graph Attention Network
GMMs Gaussian Mixture Models
LA Logical Access
LCNN Light Convolutional Neural Network
MFCCs Mel-Frequency Cepstral Coefficients
NAS Neural Architecture Search
OSI Open-Set Identification
PGD Projected Gradient Descent
ResNet Residual Networks
SiFs Speaker-irrelevant Features
SMIA Spectral Masking and Interpolation Attack
SOTA State-of-the-Art
SRS Speaker Recognition Systems
SSL Self-Supervised Learning
STFT Short-Time Fourier Transform
SV Speaker Verification
SVMs Support Vector Machines
TDNN Time Delay Neural Network
TTS Text-to-Speech
VA Voice Authentication
VAS Voice Authentication System
VC Voice Conversion
@@ 2.2. Speaker Recognition Systems (SRS)
SRS identifies individuals through their vocal features. The authentication process typically unfolds in three key phases, as shown in Figure. 1.
@@ 2.3. Anti-Spoofing Countermeasures (CMs)
Fake voice detection has evolved considerably over the past decade. Early systems primarily relied on conventional machine learning pipelines that extracted handcrafted acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and Constant-Q Cepstral Coefficients (CQCCs), which were then passed into classifiers such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) (Patel and Patil, 2015; De Leon et al., 2017). While effective at the time, these methods struggled with generalization against unseen spoofing attacks.
The field has since shifted towards deep learning methods, particularly convolutional and recurrent architectures designed to learn discriminative spectral and temporal patterns directly from audio spectrograms (Lavrentyeva et al., 2019). More recently, the state-of-the-art has been driven by Self-Supervised Learning (SSL) models used as powerful front-end feature extractors. Models such as Wav2Vec2.0 (Baevski et al., 2020) and especially WavLM (Chen et al., 2022) have become dominant, with WavLM showing superior performance due to its denoising and masked speech prediction pre-training objectives, which capture fine-grained speaker-related artifacts commonly found in deepfakes. These SSL embeddings are typically followed by smaller, task-specific back-end classifiers such as the Light Convolutional Neural Network (LCNN) (Lavrentyeva et al., 2017) or the AASIST model (Jung et al., 2022), which specialize in spoofing detection.
In parallel, another leading paradigm involves fully end-to-end systems operating directly on the raw waveform, removing the need for explicit feature engineering. The RawNet family exemplifies this approach: RawNet2 (Tak et al., 2021b) demonstrated strong performance in the ASVspoof 2019 challenge (et al., 2019), outperforming its predecessor and serving as a competitive baseline. Other end-to-end architectures have emerged as state-of-the-art alternatives. Raw-PC-DARTS (Ge et al., 2021) applies differentiable neural architecture search (NAS) to optimize network design directly on raw audio, significantly improving efficiency and detection accuracy. Meanwhile, Raw-GAT-ST (Tak et al., 2021a) integrates graph attention networks with spectro-temporal modeling to enhance the robustness of spoof detection against unseen attacks. Both approaches highlight the shift towards more adaptive and structurally optimized architectures in the fight against audio deepfakes.
@@ 2.4. Attacks Taxonomies
Attacks can be categorized based on the attacker’s goal to:
  1. (1)
@@ 2.5. Deepfake Voice Attacks
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation (Ren et al., 2019, 2020; Kim et al., 2020; Kaneko and Kameoka, 2018; Kameoka et al., 2018; Qian et al., 2019; Li et al., 2023; Qin et al., 2023; Wang et al., 2024). Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce CMs to prevent such attacks (Almutairi and Elgibreen, 2022).
One of the earliest demonstrations of AI-powered voice spoofing was conducted by Saxena et al. (Shonesy, 2015). Using a basic voice morphing tool, they were able to reconstruct a target speaker’s voice with only a few minutes of enrollment data. The spoofed voices were tested against both automated speaker verification systems and human listeners. Results revealed significant vulnerabilities: state-of-the-art speaker recognition systems failed to reject 80–90% of spoofed attempts, while human listeners correctly detected fake voices only about 50% of the time—particularly when the attacker mimicked familiar individuals.
Building on this, Alali et al. (Alali and Theodorakopoulos, 2025) explored partial voice attacks in real-world contexts. Their study demonstrated that both humans and speaker recognition systems remain vulnerable when exposed to partial speech generated through zero-shot text-to-speech and voice conversion models. The findings highlighted a high success rate in bypassing verification mechanisms and deceiving listeners. However, the study’s broad evaluation across different attacks and platforms lacked a deeper analysis of specific conditions, such as noisy environments, language variations, or system-level defenses, limiting its applicability for improving individual systems.
@@ 2.6.1. White-box Attacks
Early work by Liu et al. (Liu et al., 2019) applied white-box adversarial techniques, notably the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), to countermeasure (CM) systems. Even minimal perturbations (with $\epsilon$ below human perceptibility) caused error rates to spike dramatically, reaching 85–95% in some cases. These findings showed that, when attackers have access to model gradients, the CM’s ability to detect spoofed speech can be almost completely undermined, effectively reducing performance to random guessing.
Expanding on this, Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) introduced the Adversarial Biometrics Transformation Network (ABTN), a trainable generator that optimizes perturbations using both CM and ASV losses. Compared to FGSM, ABTN produced perturbations that generalized better across spoof types while still preserving speaker identity, thereby fooling CMs without triggering alarms in ASV systems. The method improved attack equal error rates (EER) by 27–51
In a related study, Alanis et al. (Gomez-Alanis et al., 2022b) leveraged a generative adversarial network (GAN) to deceive both CM and speaker verification (SV) systems while maintaining speaker identity. Although highly effective, this approach required access to the target or a surrogate model and concentrated solely on attack generation, without proposing defensive strategies—limiting its practical applicability.

### v3

@@ Abstract.
Voice Authentication Systems (VAS) are increasingly integrated into high-security sectors such as banking and healthcare. Despite their improvements in the previous years, they still face severe vulnerabilities from sophisticated threats, such as deepfakes and adversarial attacks. The emergence of realistic voice cloning complicates detection, as systems struggle to distinguish authentic from synthetic audio. While anti-spoofing countermeasures (CMs) exist to mitigate these risks, many rely on static detection models that can be bypassed by novel adversarial methods, leaving a critical security gap. To demonstrate this vulnerability, we propose the Spectral Masking and Interpolation Attack (SMIA), a novel method that strategically manipulates inaudible frequency regions of AI-generated audio. By altering the voice in imperceptible zones to the human ear, SMIA creates adversarial samples that sound authentic while deceiving CMs. We conducted a comprehensive evaluation of our attack against state-of-the-art (SOTA) models across multiple tasks, under simulated real-world conditions. SMIA achieved SOTA attack results, demonstrating a high success rate in bypassing combined VAS/CM systems, standalone speaker verification systems, and dedicated countermeasures. These findings conclusively show that current security postures are insufficient against adaptive adversarial attacks, highlighting the urgent need for next-generation, dynamic defenses.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM Asia Conference on Computer and Communications Security; June 01–05, 2026; Bengaluru, Indiaisbn: 978-1-4503-XXXX-X/2026/06ccs: Security and privacy Biometricsccs: Computing methodologies Neural networks
@@ 1. Introduction
Voice authentication has emerged as a widely adopted biometric technology due to its ease of use and flexibility in enhancing security within various scenarios, including mobile devices, virtual assistants, secure communication systems, financial service institutions, and law enforcement and jurisdictional departments (singh2012applications; algabri2017automatic; rose2006technical). These systems analyze an individual’s voice to create a distinct vocal signature used for identity verification.
@@
Compared to passwords or access cards, voice authentication offers reliable user verification, reduces fraud risk, and simplifies the user experience by eliminating the need to remember multiple credentials and avoiding password-related vulnerabilities. It is also more stable and cost-effective than other biometric systems. Unlike facial recognition, which requires frequent updates due to changes in appearance, voice characteristics remain relatively consistent after adulthood. Additionally, it relies on affordable microphones and requires less storage than high-resolution cameras and facial data. Additionally, it is easier to integrate with cloud services and multiple devices (renz2023authentication; springer2023twofactor; techlocity2021biometrics; khan2010comparative).
@@
Recent advancements in deep learning have revolutionized the Voice Authentication System (VAS), significantly improving accuracy, robustness, and scalability. Modern deep learning-based approaches, such as Deep Neural Network (DNN) models, have enhanced the field by creating digital voice fingerprints that serve as reference points for authentication, for example, Deep Speaker (deepspeaker) and X-Vectors (xvectors). Despite these improvements, VAS remain highly vulnerable to various forms of adversarial attacks (lan2022adversarial).
@@
In parallel, deep learning has transformed the field of synthetic speech generation, making it easier than ever to replicate someone’s voice with near-human indistinguishability. Technologies like text-to-speech (TTS) (fish-speech-v1.4; qin2023openvoice; amazon_polly) and voice conversion (VC) (li2023freevc; huang2024diffvc+) can generate highly realistic speech using minimal input data. These tools, while impressive, have raised significant concerns regarding their misuse in security-sensitive applications, particularly when synthetic speech is used to deceive or bypass VAS (shonesy2015voice; alali2025partial).
@@
To combat these deepfake attacks, various countermeasures (CMs), like RawNet2 (rawnet2), RawGAT-ST (rawgat), and RawPC-Darts (rawdarts), have been developed to distinguish between genuine and synthetic speech. However, attackers have engineered a new class of threat, known as adversarial spoofing attacks, designed specifically to deceive these CMs. This hybrid approach begins by generating a convincing spoofed audio sample using TTS or VC. Subsequently, a carefully crafted and often imperceptible adversarial perturbation is added to this synthetic speech. The resulting signal has a dual objective: bypassing the primary VAS by mimicking the target speaker while simultaneously fooling the CM into classifying the spoofed audio as authentic, posing a compounded threat to security protocols (hai2023sifdetectcracker; kassis2023breaking).
@@
A compelling example is SiFDetectCracker (hai2023sifdetectcracker), an adversarial attack that can deceive SOTA CMs by only modifying speaker-irrelevant features, such as background noise and silent segments. This method achieved an average success rate of over 80% without altering the speaker’s actual voice, demonstrating that detectors often over-rely on secondary acoustic cues. However, their impact on VAS or hybrid systems, which consist of VAS and CMs, remains unexplored. This highlights a critical gap: the need for robust attacks that are effective in real-world scenarios.
@@
Recent attacks on VAS have shifted focus from fooling speaker verification systems to breaking their CMs. Kassis et al. (kassis2023breaking) introduced a model-agnostic attack that bypasses these CMs without relying on query-based optimization, but by applying a series of universal signal-processing transformations to spoofed audio. Their approach, which manipulates cues like silence and spectral energy distribution, is capable of bypassing the entire VA stack. However, their results reveal a significant performance drop against more modern and robust defenses; for instance, the success rate fell to just 15.55% against a combined defense using the AWS Voice ID (amazon2025voiceid) and RawPC-Darts (rawdarts). This inconsistency highlights that while targeting universal flaws in CMs is a promising direction, a critical gap remains in developing attacks that can reliably defeat the current generation of SOTA, hardened VA systems.
Our main contributions are summarized as follows:
@@
  • We conduct extensive evaluations of SMIA against widely adopted SOTA open-source and commercial VAS, and SOTA CMs.
@@
Unlike prior work, which explored TTS-based attacks targeted at VAS and CMs in isolation or relied on transferability-based attacks with limited effectiveness, we present a black-box attack that simultaneously compromises both. Our evaluations span a wide range of models, tasks, and real-world scenarios, demonstrating that VAS are significantly more fragile than previously assumed, underscoring the urgent need for stronger defenses and providing a solid foundation for future research in biometric security.
This paper is organized as follows. We first review the background and related work (Section 2). We then detail our SMIA methodology (Section 3) and the experimental setup (Section 4). Section 5 presents the evaluation of our attack. Section 6 reports the results of our experiments. Section 7 provides an ablation study, analyzing the impact of different components and design choices. Finally, we discuss our findings (Section 8), outline future work (Section 9), and conclude (Section 10).
@@ 2.1. Speaker Recognition Systems (SRS)
SRS identifies individuals through their vocal features. The authentication process typically unfolds in three key phases, as shown in Figure 1:
@@ 2.2. Anti-Spoofing Countermeasures (CMs)
Fake voice detection has evolved considerably over the past decade. Early systems primarily relied on conventional machine learning pipelines that extracted handcrafted acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and Constant-Q Cepstral Coefficients (CQCCs), which were then passed into classifiers such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) (patel2015svm; de2017intro). While effective at the time, these methods struggled with generalization against unseen spoofing attacks.
The field has shifted towards deep learning methods (lavrentyeva2019stc), with SOTA now driven by Self-Supervised Learning (SSL) models like Wav2Vec2.0 (baevski2020wav2vec) and WavLM (chen2022wavlm). WavLM has shown superior performance by capturing fine-grained artifacts commonly found in deepfakes. However, pre-training these large-scale SSL models takes a lot of time and computational resources. These SSL embeddings are typically followed by smaller, task-specific back-end classifiers such as the Light Convolutional Neural Network (LCNN) (lavrentyeva2017audio) or the AASIST model (jung2022aasist), which specialize in spoofing detection.
In parallel, another leading paradigm involves fully end-to-end systems operating directly on the raw waveform, removing the need for explicit feature engineering. The RawNet family exemplifies this approach: RawNet2 (rawnet2) demonstrated strong performance in the ASVspoof 2019 challenge (Todisco2019ASVspoof), outperforming its predecessor and serving as a competitive baseline. Other end-to-end architectures have emerged as SOTA alternatives. RawPC-Darts (rawdarts) applies differentiable neural architecture search (NAS) to optimize network design directly on raw audio, significantly improving efficiency and detection accuracy. Meanwhile, RawGAT-ST (rawgat) integrates graph attention networks with spectro-temporal modeling to enhance the robustness of spoof detection against unseen attacks. Both approaches highlight the shift towards more adaptive and structurally optimized architectures in the fight against audio deepfakes.
@@ 2.3. Attacks Taxonomies
Attacks are categorized based on the attacker’s goal to:
@@
Paper Year Tasks Attack Type System Target
Deepfake Voice Attacks
Saxena et al. (shonesy2015voice) 2015 SV Black-box (Voice morphing) SV, Human listeners
Alali et al. (alali2025partial) 2025 SV Black-box (Zero-shot TTS / VC) SV, Human listeners
Adversarial Spoofing Attacks
Liu et al. (liu2019adversarial) 2019 CM White-box (FGSM/PGD) ASVspoof CM
Gomez-Alanis et al. (gomez2022adversarial) 2022 SV + CM White-box (ABTN) ASVspoof systems
Gomez-Alanis et al. (gomez2022ganba) 2022 SV + CM White-box (GANBA) ASVspoof systems
Hai et al. (hai2023sifdetectcracker) 2023 CM Black-box CM
Kassis et al. (kassis2023breaking) 2023 SV + CM Black-box (OTA spoofing) VA systems (SV + CM)
@@ 2.4. Deepfake Voice Attacks
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation (ren2019fastspeech; ren2020fastspeech; kim2020glow; kaneko2018cyclegan; kameoka2018stargan; qian2019autovc; li2023freevc; qin2023openvoice; wang2024maskgct). Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce CMs to prevent such attacks (almutairi2022review).
One of the earliest demonstrations of AI-powered voice spoofing was conducted by Saxena et al. (shonesy2015voice). Using a basic voice morphing tool, they were able to reconstruct a target speaker’s voice with only a few minutes of enrollment data. The spoofed voices were tested against both automated speaker verification systems and human listeners. Results revealed significant vulnerabilities: SOTA SRS failed to reject 80–90% of spoofed attempts, while human listeners correctly detected fake voices only about 50% of the time—particularly when the attacker mimicked familiar individuals.
Building on this, Alali et al. (alali2025partial) explored partial voice attacks in real-world contexts. Their study demonstrated that both humans and SRS remain vulnerable when exposed to partial speech generated through zero-shot TTS and VC models. The findings highlighted a high success rate in bypassing verification mechanisms and deceiving listeners. However, the study’s broad evaluation across different attacks and platforms lacked a deeper analysis of specific conditions, such as noisy environments, language variations, or system-level defenses, limiting its applicability for improving individual systems.
@@ 2.5.1. White-box Attacks
Early work by Liu et al. (liu2019adversarial) applied white-box adversarial techniques, notably the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), to CM systems. Even minimal perturbations (with $\epsilon$ below human perceptibility) caused error rates to spike dramatically, reaching 85–95% in some cases. These findings showed that, when attackers have access to model gradients, the CM’s ability to detect spoofed speech can be almost completely undermined, effectively reducing performance to random guessing.
Expanding on this, Gomez-Alanis et al. (gomez2022adversarial) introduced the Adversarial Biometrics Transformation Network (ABTN), a trainable generator that optimizes perturbations using both CM and ASV losses. Compared to FGSM, ABTN produced perturbations that generalized better across spoof types while still preserving speaker identity, thereby fooling CMs without triggering alarms in ASV systems. The method improved attack equal error rates (EER) by 27–51% (absolute) across ASVspoof 2019 logical and physical access protocols (Todisco2019ASVspoof).
In a related study, Alanis et al. (gomez2022ganba) leveraged a generative adversarial network (GAN) to deceive both CM and speaker verification (SV) systems while maintaining speaker identity. Although highly effective, this approach required access to the target or a surrogate model and concentrated solely on attack generation, without proposing defensive strategies—limiting its practical applicability.